# Video LLMs for Temporal Reasoning in Long Videos

### v4

@@ 3.2 Bidirectional Long Short-Term Memory
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$. Please refer to the textbook [19] for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As empirically observed in Sec. 5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [44]. Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike LongVLM [50], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec. 5.
The video tokens output by our BiLSTM module have dimensions of $(C*N^{c}_{f},~{}2*N_{V})$, while the LLM requires an input of dimensions $(N^{c}_{f},~{}N_{LLM})$. Here, $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM. Thus, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
@@ 3.3 Large Language Model
The LLM (i.e., yellow block in Fig. 2) takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users. Video LLMs typically employ a two-stage training scheme: i) leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and ii) utilizing instruction data for instruction following to fine-tune the pre-trained model. Following TimeChat [36], we use the checkpoint of the open-source LLaMA-2 7B model [43] and perform instruction tuning only. We optimize the below loss function:
@@
Here, $\mathbf{\psi}$ is the learnable parameters of TemporalVLM, $L$ is the response length, $x_{i}$ is the current predicted token, and $\mathbf{X}^{r}_{<i}$ is the prior tokens appearing before $x_{i}$ in the response. Also, we employ LoRA [12] for efficient fine-tuning.
4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
We introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies.
@@ 4.1 Dataset Statistics
Our IndustryASM dataset includes 4803 videos with each video lasting 119 seconds on average, yielding 158 hours as the total dataset duration. In addition, it consists of 52 various industry assembly procedures with each assembly procedure having 12 steps on average. Example videos are shown in Fig. 3. Due to space limits, we provide additional details of IndustryASM in our supplementary material.
@@ 4.2 Annotation Process
Our industry assembly videos are manually annotated by industrial engineers into relevant framewise action segments. The action labels contain the same step naming conventions as used on factory floors during manufacturing processes. To ensure quality, each video is annotated by two labelers, i.e., one labeler obtains the labels for the video, while another labeler checks the labels. In cases of conflicts, both labelers discuss to address them. Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of roughly 92% between labelers. Given the framewise annotations we first convert them to timestamps by using the frame rate of the video. We manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using timestamps and action labels. Example instructions and responses are shown in Fig. S1 of our supplementary material.
@@ 5 Experiments
Tasks and Datasets. Our TemporalVLM model is evaluated on a few temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. For fair comparisons with TimeChat [36], we fine-tune our model on a subset [^1] We could not download the YT-Temporal [55] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [36] and Valley [28] datasets. The combined dataset contains a total of 142K long videos with diverse activities. We follow TimeChat [36] to use 6 instructions (generated by GPT-4 [1]) per task and convert the video annotations into chat samples. We evaluate on YouCook2 [59] for dense video captioning, Charades-STA [8] for temporal video grounding, QVHighlights [20] for video highlight detection, and our IndustryASM dataset for temporal action segmentation.
@@
Implementation Details. We follow TimeChat [36] to use LLaMA-2 7B [43] as the LLM, and ViT-G/14 from EVA-CLIP [40] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [57] checkpoints to initialize the video Q-Former. We employ pytorch’s [15] implementation of multiheaded attention in the fusion module of the clip encoder with 8 attention heads. The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [12] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [36] and Valley [28] on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout fine-tuning. Please see our supplementary material for more details.
@@
Competing Methods. We compare our model against state-of-the-art video LLMs, including VideoChat-Embed [22], Video-LLaMA [57], Valley [28], TimeChat [36], and LongVLM [50]. Since LongVLM [50] is not trained on TimeIT [36] and Valley [28] and it is a multi-model pipeline (TimeChat [36] and TemporalVLM are end-to-end models), we replace our time-aware clip encoder and BiLSTM modules with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM model which we use for evaluation. Nevertheless, we compare with the original LongVLM model [50] on general video understanding in our supplementary material.
Evaluation Metrics. We utilize different sets of metrics for evaluating temporal reasoning tasks:
@@
  • Dense video captioning: The overall story-level evaluation is done through SODA_c [7]. The event localization performance is measured using F1 Score. CIDEr [45] is used to measure the quality of the generated captions.
  • Temporal video grounding: To gauge the effectiveness of temporal video grounding, we use R@1(IoU = $\alpha$), which is the percentage of predictions with an intersection over union greater than $\alpha$ compared to the ground truth.
@@
  • Video highlight detection: We use mean average precision (mAP) and HIT@1 at IoU= 0.75 to evaluate the accuracy of the saliency scores for key moments in videos.
@@
  • Temporal action segmentation: We report the framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance. We use the intersection over union ratio to find the overlapping thresholds.
@@ 5.1 Zero-Shot Results
Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
YouCook2 [59] QVHighlights [20] Charades-STA [8]
@@
Valley [28] 0.1 0.0 1.5 10.9 15.2 4.7 1.6
Video-LLaMA [57] 0.0 0.0 0.1 11.3 15.6 2.7 1.2
VideoChat-Embed [22] 0.2 0.6 3.4 13.1 18.1 3.2 1.4
LongVLM [50] 0.8 2.5 12.3 11.1 15.0 13.9 6.1
TimeChat [36] 1.2 3.4 12.6 14.5 23.9 27.9 12.3
TemporalVLM (Ours) 1.2 3.7 13.1 16.4 31.3 30.1 13.2
Table 1: Zero-short results. We evaluate all methods on YouCook2 [59] for dense video captioning, QVHighlights [20] for video highlight detection, and Charades-STA [8] for temporal video grounding. Best results are in bold, while second best ones are underlined.
Dense Video Captioning Results. Tab. 1 shows that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score. For example, we outperform TimeChat [36] by +0.3 on CIDEr and +0.5 on F1 Score. This is likely because TimeChat [36] treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information. In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization. Moreover, our approach obtains superior results than LongVLM [50] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
Video Highlight Detection Results. The video highlight detection results in Tab. 1 show that our TemporalVLM model obtains the best performance on both mAP and HIT@1. In particular, it surpasses TimeChat [36] by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM [50] by +5.3 at mAP and by +16.3 at HIT@1. These results suggest the effectiveness of our model for video highlight detection. By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their ranks of importance within the video against a given query. Though LongVLM [50] extracts multi-level features, the lack of a time-aware encoder hinders their results.
Temporal Video Grounding Results. It is evident from Tab. 1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [36] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [50] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM module to capture long-range temporal relationships between multiple clips to detect the given event effectively. The lack of a time-aware encoder affects LongVLM [50], while TimeChat [36] suffers from a lack of fine-grained information.
@@ 5.2 Supervised Results
Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results. We present the results of all methods in the supervised setting in Tab. 2. Our TemporalVLM model achieves the best results across all tasks and metrics. As compared to the zero-shot results in Tab. 1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task. However, we acknowledge that task-specific models such as MMN [49] and QD-DETR [30] may outperform generalist models such as TimeChat [36] and our TemporalVLM model due to their task-specific design paradigms, e.g., using a mutually matching network to directly model the relationship between the query text and video content or specialized tokens and loss functions during fine-tuning.
Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
YouCook2 [59] QVHighlights [20] Charades-STA [8]
@@
LongVLM [50] 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat [36] 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM (Ours) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 2: Supervised results. We first fine-tune all models on TimeIT [36] and Valley [28], and then perform task-specific fine-tuning and evaluation on YouCook2 [59] for dense video captioning, QVHighlights [20] for video highlight detection, and Charades-STA [8] for temporal video grounding. Best results are in bold, while second best ones are underlined.
Temporal Action Segmentation Results. We further test all methods on our IndustryASM dataset for temporal action segmentation. Since the activities in IndustryASM significantly differ from those in TimeIT [36] and Valley [28], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat [36] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [50] by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our fused time-aware and multi-level features.
@@
TimeChat [36] {11.1, 8.5, 4.4} 32.7 46.0
LongVLM [50] {17.1, 13.3, 7.3} 47.3 47.6
TemporalVLM (Ours) {22.3, 18.3, 11.1} 56.2 52.9
Table 3: Temporal action segmentation results in the supervised setting on IndustryASM. All models are fine-tuned on IndustryASM before evaluation. Best results are in bold, while second best ones are underlined.
Multi-Clip Encoder vs. Time-Aware Encoder. Multi-clip encoders in LongVLM [50] and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat [36] on temporal action segmentation in Tab. 3. In contrast, time-aware encoders in TimeChat [36] and TemporalVLM are important to event localization and temporal modeling, resulting in better performance than LongVLM [50] on dense video captioning, video highlight detection, and temporal video grounding in Tabs. 1 and 2. Our TemporalVLM encoder is both multi-clip and time-aware, yielding the best performance across all tasks.

### v5

@@ 3.2 Long-Term Temporal Reasoning
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward and backward LSTMs respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward and backward LSTMs as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$. Please refer to the textbook Learning (2016) for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As observed in Tab. 5, BiLSTM outperforms various alternatives, including average pooling, linear layer, LSTM, and transformer Vaswani et al. (2017). Unlike LongVLM Weng et al. (2024), which applies pooling to obtain global features, our BiLSTM module has learnable parameters and achieves better results. Advanced recurring models, e.g., state space Gu et al. (2021) and Mamba Gu and Dao (2024), may further improve the performance, which remains our future work.
The video tokens output by our BiLSTM module are of size $(C\times N^{c}_{f},~2\times N_{V})$, while the LLM requires an input size of $(N^{c}_{f},~N_{LLM})$. $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM. Thus, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
@@ 3.3 Large Language Model
The LLM takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users. Video LLMs typically employ a two-stage training: leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and utilizing instruction data to fine-tune the pre-trained model. Following TimeChat Ren et al. (2024), we use the checkpoint of the LLaMA-2 7B model Touvron et al. (2023b) and perform instruction tuning only. We optimize the below loss:
@@
Here, $\mathbf{\psi}$ is the learnable parameters of TemporalVLM, $L$ is the response length, $x_{i}$ is the current predicted token, and $\mathbf{X}^{r}_{<i}$ is the prior tokens appearing before $x_{i}$ in the response.
4 IndustryASM
Figure 3: Example IndustryASM videos with different camera viewpoints, actors, backgrounds, and activities.
@@ 4.1 Dataset Statistics
IndustryASM includes 4803 videos with each lasting 119 seconds on average, yielding 158 hours as the total dataset duration. In addition, it consists of 52 various industry assembly procedures with each having 12 steps on average. Example videos are shown in Fig. 3. Due to space limits, we provide additional details in our supplementary material.
@@ 4.2 Annotation Process
Our videos are annotated by industrial engineers into framewise action segments. The action labels contain the same step naming conventions as used on factory floors during manufacturing processes. To ensure quality, each video is annotated by two labelers, i.e., one obtains the labels for the video, while another checks the labels. In cases of conflicts, both labelers discuss to address them. Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of 92% between labelers. Given the framewise annotations we first convert them to timestamps by using the frame rate of the video. We then manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using the timestamps and action labels. Examples are included in our supplementary material.
@@ 5 Experiments
Tasks and Datasets. We evaluate TemporalVLM on dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. For fair comparisons with TimeChat Ren et al. (2024), we fine-tune our model on a subset [^1] We could not download YT-Temporal Zellers et al. (2022) due to its large size and the restricted number of downloads. of TimeIT Ren et al. (2024) and Valley Luo et al. (2023b). The combined dataset contains 142K long videos with diverse activities. We follow TimeChat to use 6 instructions (generated by GPT-4 Achiam et al. (2023)) per task and convert video annotations into chat samples. We evaluate on YouCook2 Zhou et al. (2018) for dense video captioning, Charades-STA Gao et al. (2017) for temporal video grounding, QVHighlights Lei et al. (2021) for video highlight detection, and IndustryASM for temporal action segmentation. Note that only Tab. 3 uses our dataset, all the remaining evaluations use existing datasets. Following prior works, we report the results from a single run.
@@
Implementation Details. We follow TimeChat Ren et al. (2024) to use LLaMA-2 7B Touvron et al. (2023b) as the LLM, and ViT-G/14 from EVA-CLIP Sun et al. (2023) as the image encoder. We take InstructBLIP’s Dai et al. (2023) checkpoints to initialize the image Q-Former, and Video-LLaMA’s Zhang et al. (2023) checkpoints to initialize the video Q-Former. We employ pyTorch’s Imambi et al. (2021) multiheaded attention in the fusion module. The BiLSTM module includes forward and backward LSTMs with one hidden layer each. The BiLSTM and projection layers are initialized randomly, while the LLM and image encoder remain frozen throughout fine-tuning. The LLM is fine-tuned using LoRA Hu et al. (2023). We fine-tune our model on an 8-A100 (80 GB) machine. Please see our supplementary material for more details.
@@
Competing Methods. We compare our model against VideoChat-Embed Li et al. (2023b), Video-LLaMA Zhang et al. (2023), Valley Luo et al. (2023b), TimeChat Ren et al. (2024), and LongVLM Weng et al. (2024). Since LongVLM is not trained on TimeIT Ren et al. (2024) and Valley Luo et al. (2023b) and it is a multi-model pipeline (TimeChat and TemporalVLM are end-to-end), we replace our time-aware clip encoder and BiLSTM with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM for evaluation. Nevertheless, we compare with the original LongVLM on general video understanding in our supplementary material.
Evaluation Metrics. We use the below metrics:
@@
  • Dense video captioning: We use SODA_c Fujita et al. (2020) for story evaluation, F1 Score for event localization, and CIDEr Vedantam et al. (2015) for measuring caption quality.
[…]
@@
  • Video highlight detection: We use mean average precision (mAP) and HIT@1 (IoU= 0.75) to evaluate saliency scores of key moments.
@@
  • Temporal action segmentation: We use framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance.
@@ 5.1 Zero-Shot Results
Model Dense Captioning Highlight Detection Temporal Grounding
YouCook2 QVHighlights Charades-STA
@@
Valley Luo et al. (2023b) 0.1 0.0 1.5 10.9 15.2 4.7 1.6
Video-LLaMA Zhang et al. (2023) 0.0 0.0 0.1 11.3 15.6 2.7 1.2
VideoChat-Embed Li et al. (2023b) 0.2 0.6 3.4 13.1 18.1 3.2 1.4
LongVLM Weng et al. (2024) 0.8 2.5 12.3 11.1 15.0 13.9 6.1
TimeChat  Ren et al. (2024) 1.2 3.4 12.6 14.5 23.9 27.9 12.3
TemporalVLM 1.2 3.7 13.1 16.4 31.3 30.1 13.2
Table 1: Zero-short results. We test on YouCook2 for dense video captioning, QVHighlights for video highlight detection, Charades-STA for temporal video grounding. Best results are in bold, second best ones are underlined.
Dense Video Captioning Results. Tab. 1 shows that TemporalVLM achieves the best results across all metrics. For example, we outperform TimeChat Ren et al. (2024) by +0.3 on CIDEr and +0.5 on F1 Score. This is likely because TimeChat treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information. In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization. Our approach also obtains superior results than LongVLM Weng et al. (2024) across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
Video Highlight Detection Results. From Tab. 1, TemporalVLM obtains the best performance on both mAP and HIT@1. It surpasses TimeChat Ren et al. (2024) by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM Weng et al. (2024) by +5.3 at mAP and by +16.3 at HIT@1. These results suggest the effectiveness of our model for video highlight detection. By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their relative importance within the video for a given query. Though LongVLM extracts multi-level features, the lack of a time-aware encoder hinders its results.
Temporal Video Grounding Results. It is evident from Tab. 1 that TemporalVLM achieves the best performance across R@1(IoU=0.5) and R@1(IoU=0.7). It outperforms TimeChat Ren et al. (2024) by +2.2 and +0.9 on R@1(IoU=0.5) and R@1(IoU=0.7) respectively, and LongVLM Weng et al. (2024) by +16.2 and +7.1 on R@1(IoU=0.5) and R@1(IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM to capture long-range temporal relationships across clips to detect the given event effectively. The lack of a time-aware encoder affects LongVLM, while TimeChat suffers from a lack of fine-grained cues.
@@ 5.2 Supervised Results
Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results. Tab. 2 presents the results in the supervised setting. TemporalVLM performs the best across all tasks and metrics. As compared to the zero-shot results in Tab. 1, we see an increase in performance, e.g., +24.3 on R@1(IoU=0.5) in temporal video grounding and +9.5 on CIDEr in dense video captioning. However, we acknowledge that task-specific models, e.g., MMN Wang et al. (2022) and QD-DETR Moon et al. (2023), may outperform generalist models, e.g., TimeChat Ren et al. (2024) and TemporalVLM, due to their task-specific designs. Due to space limits, we provide a comparison with task-specific models in our supplementary material.
Model Dense Captioning Highlight Detection Temporal Grounding
YouCook2 QVHighlights Charades-STA
@@
LongVLM Weng et al. (2024) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat Ren et al. (2024) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 2: Supervised results. We first fine-tune all models on TimeIT and Valley, and then perform task-specific fine-tuning and evaluation on YouCook2 for dense video captioning, QVHighlights for video highlight detection, and Charades-STA for temporal video grounding. Best results are in bold, while second best ones are underlined.
Temporal Action Segmentation Results. We now evaluate on our IndustryASM dataset. Since the activities in IndustryASM significantly differ from those in TimeIT Ren et al. (2024) and Valley Luo et al. (2023b), we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat Ren et al. (2024) by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM Weng et al. (2024) by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our time-aware and multi-level features.
@@
LongVLM Weng et al. (2024) {17.1, 13.3, 7.3} 47.3 47.6
TimeChat Ren et al. (2024) {11.1, 8.5, 4.4} 32.7 46.0
TemporalVLM {22.3, 18.3, 11.1} 56.2 52.9
Table 3: Temporal action segmentation results in supervised setting on IndustryASM. All models are fine-tuned on IndustryASM before evaluation. Best results are in bold, while second best ones are underlined.
Multi-Clip Encoder vs. Time-Aware Encoder. Multi-clip encoders in LongVLM Weng et al. (2024) and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat Ren et al. (2024) on temporal action segmentation in Tab. 3. In contrast, time-aware encoders in TimeChat and TemporalVLM are important to event localization and temporal modeling, yielding better results than LongVLM on dense video captioning, video highlight detection, and temporal video grounding in Tabs. 1 and 2. TemporalVLM is both multi-clip and time-aware.